An electron is to be accelerated form 3.4 × 106 m/s at point A to 8.3 × 106 m/s at point B. Through what potential difference (VB- VA) must the electron pass to accomplish this? (Give your answer in decimal using V as unit)
Answer #1

change in Kinetic energy = q(Vb - Va)

0.5mvb^2 - 0.5mva^2 = q(Vb - Va)

0.5*9.11*10^-31*((8.3*10^6)^2 - (3.4*10^6)^2 ) = 1.6*10^-19*(Vb - Va)

(Vb - Va) = 163.2 V

so the potential difference is 163.2 V